A man cleaning his apartment pulls the canister of a vacuum cleaner with a force of magnitude F = 50.0 N at an angle of 30.0°. He moves the vacuum cleaner a distance of 30.0 m. Calculate

i) the work done by 50.0 N force
ii) the work done on the vacuum cleaner by the man if he pulls it 3.00 m with a horizontal force of F = 50.0 N
iii) the nett work done of the system for problem (i)
1
Expert's answer
2020-10-27T11:15:31-0400

i)


"W=(50)(30)\cos{30}=1300\ J"

ii)


"W=(50)(3)=150\ J"

iii)


"W=(50)(30)\cos{30}=1300\ J"